The officer's of the science club are planning to sell 125 tickets to be raffled during the school's foundation day . One ticket will win PHP 2,000 and the other tickets win nothing. If you buy one ticket, what will be your expected gain?




1
Expert's answer
2022-02-14T16:23:25-0500

"E(X)=2000*\frac{1}{125}+0*(1-\frac{1}{125})=16."